Sol Summary – October 14th

Report title: Sol 2 Summary Report

Crew #: 319

Position: Mission Commander

Report prepared by: Rym Y. Chaid

Date: 14-10-2025

Sol: 2

Summary Title: Canadian Turkey spotted on Mars!

Mission Status: Nominal. All systems operational.

Sol Activity Summary:

This is Commander Rym Y. Chaid reporting from Mars Station.

SOL 2 reminded us that life on Mars, even an analog Mars, is never predictable. Early in the day, one of our GreenHab experiments ran into technical difficulties during setup. What was expected to be a simple procedure quickly became a full-crew operation, and all engineers on deck responded to action. Ricardo, Shriya and myself, all came in to troubleshoot and support our GreenHab Officer Mackenzie in this difficult task.

In moments like these, you see the strength of a crew, not in how perfectly things go, but in how quickly everyone steps in when they don’t. There’s no “someone else’s task” here on Mars. Every system depends on every person, and today, we proved just that.

On a happy note, the habitat was filled with anticipation all day as we’re preparing for our first EVA tomorrow. The suits are ready, checklists complete, and spirits high. You can feel the excitement in every conversation, every glance toward the airlock.

And since I am proudly Canadian, we’re closing SOL 2 by celebrating Canadian Thanksgiving on Mars, proof that even 225 million kilometers from home, you can still find something to be grateful for… especially if someone remembered the instant mashed potatoes!

As we wrap up SOL 2, there’s a sense of calm pride in the air. We adapted, we worked together, and we grew stronger as a team. Each Sol brings its own story, and today it was about teamwork, flexibility, and trust.

Tomorrow, we take our first steps outside, on the surface of Mars, a moment every astronaut dreams of.

End Transmission.

Look Ahead Plan: First Crew EVA tomorrow, pending approval and weather conditions.

Anomalies in work: One GreenHab experiment had technical difficulties during setup. All engineers on deck had to step up: Commander, Scientist and Engineers came in as support for GreenHab Officer.

Weather: Temperature around 57 F, No rain, 67% Humidity. Overall Sunny day with slight overcast

Crew Physical Status: All Crew in Good Health + Spirit.

EVA: No EVA activity today. Request sent it for tomorrow – SOL 3.

Reports to be filed: Sol Summary, Journalist Report, Crew Photos, Greenhab Report, Operations Report, Mission Plan

Support Requested: None at this time.

GreenHab Report – October 17th

Green Hab Report

Crew #319
Position: Crew Journalist & Green Hab Officer
Report prepared by: Mackenzie Calle
Date: 17-10-2025
Sol: 5

Environmental control (fan & heater): Fan min 75F, max 80F
Average temperatures (last 24h): 86.6F
Maximum temperature (last 24h): 124.5F
Minimum temperature (last 24h): 48.7F
Hours of supplemental light: 0
Daily water usage for crops: 0.1 gallons
Daily water usage for research and/or other purposes: 0
Water in the Tank (160 gal useful capacity): _160____ gallons remaining (system says 160 but I used some last night and the last few days it has been at 159)
Time(s) of watering for crops: 5 minutes misting the microgreen seeds for Spaceseed and MDRS microgreens
Changes to crops: Nothing to report
Soil Moisture Level: 9.4 (Average of five points, one on each corner and center)

Green Hab Officer checked on both Spaceseed and the microgreens experiments, both of which are currently in trays that are stacked and in dark locations with weight on top of them to encourage germination. Spaceseed has started to germinate evenly across the top tray but the bottom three trays germination has only begun in locations where the ridges of the tray on top does not press into the seeds as strongly.

The microgreens with moist paper towels have not started germination and have again been watered. There is some concern over the 124.5F max temperature that was reached today and the seeds not having enough water to begin germination, especially for the paper towels. The seeds with vermiculite that was soaked in water have just begun some germination but the seeds where vermiculite was moist have not.

Journalist Report – October 17th

Journalist Report
Crew 319
Position: Crew Journalist
Report prepared by: Mackenzie Calle
Date: 17-10-2025
Sol: 5

Report from Sol 5

Reporting from Sol 5 10.17.25 5:59PM. The theme of the day was two – two EVAs, with two crew each, for 2.5 hours each. Crew Scientist Shriya Musuku and Crew Journalist Mackenzie Calle embarked on the first EVA this morning, departing in Spirit Rover. Crew was inspired by the landscape as this was the farthest we have ventured since being on Mars. After passing Galileo Road however, crew lost comms with Habcom and spent most of the EVA trying to reestablish communication and understand where we were losing signal. Crew eventually needed to abort mission and make their way back to the hab.

EVA Crew 2 today with Commander Rym Chaid and Crew Engineer Ricardo J. Gonzalez found significantly more success with radio communication. We established that the comms were significantly better when facing in the direction of North Ridge. Crew made their way to a different destination than intended but walked to the beautiful North Ridge hills to begin geological sampling.

Overall crew is feeling good about the EVAs and learned a lot of lessons that we can carry forward for the rest of the mission.

Operations Report – October 17th

Report title: Operations Report
Crew #: 319
Position: Crew Engineer
Report prepared by: Ricardo Javier Gonzalez
Date: 17-10-2025
Sol: 5

NON-NOMINAL SYSTEMS

Non-nominal systems:
Nothing to report

Notes on non-nominal systems:
Nothing to report

ROVERS

Spirit rover used: Yes
Hours:
Initial: 283.8
Final: 284.4
Beginning charge: 100%
Ending charge: 66%
Currently charging: No

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: Yes
Hours:
Initial: 370.1
Final: 372.30
Beginning charge: 100%
Ending charge: 70%
Currently charging: No

General notes on rovers:
All rovers nominal post EVA – nothing to report.

CAMPUS

Summary of Hab operations:
Water Use (see notes): 7.98 gal/d
Main tank level (remaining gallons, see notes): 290 gal (69.6%)
Static tank pipe heater (ON or OFF) (Orange cable): ON
Static tank heater (ON or OFF) (Black cable): ON
Toilet tank emptied (NO or YES): No

Summary of internet:
Nothing to report

Summary of suits and radios:
Suits – all suits are nominal post EVA.

Radios – Crew engineer will attempt to re-attach the broken earpiece to head set #1 using glue and tape. If unsuccessful, crew engineer will leave headset in RAM for Mission Support.

All other radios are nominal.

Summary of GreenHab operations:
Water use: 0.1 gal
Heater (ON or OFF): ON
Supplemental light (hours of operation): 0 hrs
Harvest (name, weight in grams): N/A

Summary of ScienceDome operations:
Nothing to report
Dual split (Heat or AC, On or Off or Automatic): OFF

Summary of RAM operations:
Nothing to Report

Summary of any observatory issues:
Nothing to Report

Summary of health and safety issues:
HSO reported all crew safe and verified all fire extinguishers on stations to be nominal

Questions, concerns and requests to Mission Support:
N/A

Astronomy Report – October 16th

[category  astronomy-report]

Report title: Astronomy Report
Crew #: 319
Position: Crew Engineer
Report prepared by: Ricardo Javier Gonzalez
Date: 16-10-2025
Sol: 4

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one MDRS-14 or MDRS-WF): N/A
Objects to be Imaged this Evening: N/A
Images submitted with this report: N/A
Problems Encountered: N/A

MUSK OBSERVATORY

Solar Features Observed: Crew Engineer observed several sunspots on both the northern and southern hemispheres.
Images submitted with this report: Yes (see images below)
Problems Encountered: Crew Engineer encountered no issues during observatory operations, but had some difficulties tuning the image of the solar prominences on the edge of the Sun. A re-attempt will be made either tomorrow or the following day to adjust settings and produce a higher quality image!

Journalist Report – October 16th

Journalist Report
Crew 319
Position: Crew Journalist
Report prepared by: Mackenzie Calle
Date: 16-10-2025
Sol: 4

Report from Sol 4

I want to start off today’s report with a quote that perfectly encapsulates the day from Crew Engineer Ricardo J. Gonzalez, “We explored from the micro to the macro and looked at everything from the cells to the sun.”

Crew Scientist Shriya Musuku spent the morning in the Science Dome, working on a lot of the equipment, including our microscope! We did a few sample collections around the base and were able to look at the Martian regolith at 100x magnification. The crew is now incredibly excited about what we will be able to sample for our geology experiments, which will begin on tomorrow’s EVA.

Crew Engineer Ricardo successfully began operation of the Solar Observatory and the crew is thrilled to have this space (forgive the pun). We all took turns looking at the sun through the telescope and visualized the cool spots, also known as sunspots. We’re looking forward to imaging solar flares and combining that with the photos of the cool spots to create a full image of the sun through the telescope software.

Spaceseed is now fully operational and we are now on Sol 2 of this experiment. The microgreens experiment that compares vermiculite and a paper towel as a base for broccoli, radish, mung bean, and alfalfa sprouts. Stay tuned for our results.

Crew also spent time on media today with both Commander Rym Chaid and Crew Journalist Mackenzie Calle documenting different aspects of our mission and the base through photo and video.

All crew are really looking forward to our EVA tomorrow morning and being able to freely explore more of the Martian terrain!

GreenHab Report – October 16th

Green Hab Report

Crew #319
Position: Crew Journalist & Green Hab Officer
Report prepared by: Mackenzie Calle
Date: 16-10-2025
Sol: 4

Environmental control (fan & heater): Fan min 75F, max 80F
Average temperatures (last 24h): 68.2F
Maximum temperature (last 24h): 86F
Minimum temperature (last 24h): 50.4F
Hours of supplemental light: 0
Daily water usage for crops: 0.25 gallons (approx)
Daily water usage for research and/or other purposes: 0
Water in the Tank (160 gal useful capacity): _160____ gallons remaining (system says 160 but I used some last night and the last few days it has been at 159)
Time(s) of watering for crops: 2 hours on setting up microgreens in appropriate growth trays for MDRS experiment and Spaceseed
Changes to crops: Nothing to report
Soil Moisture Level: 9.4 (Average of five points, one on each corner and center)

Green Hab Officer completed Spaceseed setup evening of Sol 3. The structure was completed earlier in the evening and then the final technology and software was completed prior to installing the microgreens on the rockwool trays. We believe the system is properly connecting to the PIs but it did need to be unplugged and re-plugged into a new power source. From our end everything is nominal but awaiting confirmation.

The microgreens research experiment has also begun comparing vermiculite to a wet paper towel base. Broccoli, radish, mung bean, and alfalfa sprouts have all been planted and we will monitor growth over the next 8-10 days.

Sol Summary – October 16th

Report title: Sol 4 Summary Report
Crew #: 319
Position: Mission Commander
Report prepared by: Rym Y. Chaid
Date: 16-10-2025
Sol: 4

Summary Title: Beauty from Above

Mission Status: Nominal. All systems operational.

Sol Activity Summary:

This is Commander Rym Y. Chaid reporting from Mars Station.

Today, SOL 4 brought light, in every sense of the word.
Our crew woke up with renewed energy and optimism, with a shared sense of purpose. The habitat felt alive again.

Our Crew Astronomer, Ricardo, began the morning by setting up the Solar Observatory. For the first time during this mission, we turned our eyes toward our very own star, the same Sun that shines over Earth, now seen from the perspective of Mars. Through the telescope, we observed sunspots across the solar surface, standing as a reminder of the vast and dynamic universe we are part of. Standing together around the observatory dome, all of us shared a feeling of awe mixed with scientific curiosity, and the desire for exploration.

Later in the day, Crew Journalist Mackenzie initiated our first drone operations, launching one of our two aerial systems to survey the terrain surrounding MDRS . The drone flew above the red landscape, capturing an aerial view of our station. This was a breathtaking image of human presence in isolation, a small beacon of exploration against the endless expanse. For me, seeing our habitat from above felt symbolic: it was a reminder that perspective, even in small ways, can change everything.

The mood in the Hab was high and harmonious. We’ve moved beyond yesterday’s setbacks stronger, more synchronized, and ready for what lies ahead. Tomorrow marks the beginning of a new series of EVAs, and excitement runs high through every conversation and preparation checklist.

As the Commander and Mission Specialist, I spent part of the day planning our first set of long-range EVAs that will take our crew back into the vast Martian terrain. Every EVA requires precision, from navigation routes to rover coordination, terrain risk assessment, comms planning, and emergency contingencies. My goal is simple but vital: ensure that my crew is prepared, trained, and safe as we step further into the unknown. Mars doesn’t forgive mistakes, so every line on a map, every procedure rehearsed inside the Hab, and every checklist double-checked is our shield against the unforgiving beauty of this planet.

Tonight, we also celebrated a milestone beyond our own walls. A new analog astronaut habitat, the Slowak Cosmic Region Habitat, officially joined the World’s Biggest Analog Mission. At their special request, our crew recorded a welcome transmission from the Mars Desert Research Station, an honor that filled us with pride. To welcome fellow explorers into this global mission of discovery reminds us that exploration is not a solitary act, but a shared human journey.

From one habitat to another, from one world to the next, we stand united in curiosity, courage, and wonder.

Commander Rym Y. Chaid
Advanced Mars Crew

End Transmission.

Anomalies in work: None.

Crew Physical Status: All Crew in Good Health – High Spirits.

EVA: None.

Reports to be filed: Sol Summary, Journalist Report, Crew Photos, Greenhab Report, Operations Report, Mission Plan, EVA Report 1 & 2

Support Requested: None at this time.

Copyright © The Mars Society. All rights reserved. | Main Site